# Back to the phase space: thermal axion dark radiation via couplings to standard model fermions

### v1

@@ III Back to the phase space
$\begin{split}&\frac{d\rho_{\mathcal{B}}}{dt}+3H(1+w_{\mathcal{B}})\rho_{\mathcal{B}}=\\ &-g_{a}\int\frac{d^{3}k}{(2\pi)^{3}}k\,\left(1-\frac{{\mathcal{F}}_{a}(k,t)}{{\mathcal{F}}_{a}^{\rm eq}(k,t)}\right)\mathcal{C}_{a}(k,t)\ .\end{split}$ (6)
Here, $w_{\mathcal{B}}$ is the equation of state parameter [48]. Finally, the Hubble rate is given by the Friedmann equation
$H=\frac{1}{\sqrt{3}M_{\rm Pl}}\sqrt{\rho_{\mathcal{B}}+\rho_{a}}\ .$ (7)
@@ III.2 Axion temperature
Here, $\zeta(x)$ is the Riemann function evaluated for the argument $x$, and the presence of this overall factor ensures that we recover the temperature of a Bose-Einstein distribution when the interactions are strong enough to ensure thermalization. We will show how the axion temperature defined in the above equation evolves in the scenarios considered in the next two sections.
@@ IV Leptophilic Axions
If one chooses this clever set of variables then it is possible to investigate compactly in a single plot all deviations from thermal shapes for different choices of the coupling strengths. This is what the left panels of Fig. 2 illustrate. Axions produced from scatterings of charged leptons have a momentum distribution always very close to a thermal shape identified by the black dotted line. This is true for large coupling where thermalization is achieved at some point in the past as well as for small couplings where axions do not get even close to equilibrium. The fact that the asymptotic distribution is always rather thermal up to an overall constant (the factor $R_{a}(A_{F})$ in the normalization) makes sense given the thermal origin. Crucially, the resulting temperatures are different as shown by the right panels of Fig. 2. Larger values of $f_{a}$ lead to smaller temperatures, and the net observable effect is a reduced contribution to $\Delta N_{\rm eff}$ that we evaluate in Sec. VI.
@@ Acknowledgements.
The authors thank Miguel Escudero, Rotem Ovadia and Luca Vecchi for useful discussions. This work is supported in part by the Italian MUR Departments of Excellence grant 2023-2027 “Quantum Frontiers”. F.D. is supported by Istituto Nazionale di Fisica Nucleare (INFN) through the Theoretical Astroparticle Physics (TAsP) project. F.D. acknowledges support from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN. The work of A.L. is supported by an ERC STG grant (“Light-Dark”, grant No. 101040019). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement No. 101040019). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. The European Union cannot be held responsible for them. A.L. is also in part supported by the US-Israeli BSF grant 2016153. A.L. is grateful to the CERN theory group for the support while this work was carried out.
@@ Appendix B Tracking the axion number density
$n_{a}(t)\equiv\int\frac{d^{3}k}{(2\pi)^{3}}{\mathcal{F}}_{a}(k,t)\ .$ (20)

### v2

@@ III Back to the phase space
$\begin{split}&\frac{d\rho_{\mathcal{B}}}{dt}+3H(1+w_{\mathcal{B}})\rho_{\mathcal{B}}=\\ &-g_{a}\int\frac{d^{3}k}{(2\pi)^{3}}k\,\left(1-\frac{{\mathcal{F}}_{a}(k,t)}{{\mathcal{F}}_{a}^{\rm eq}(k,t)}\right)\mathcal{C}_{a}(k,t)\ ,\end{split}$ (6)
with $g_{a}=1$ and $w_{\mathcal{B}}$ the equation of state parameter [48]. The Hubble rate is given by the Friedmann equation
$H=\frac{1}{\sqrt{3}M_{\rm Pl}}\sqrt{\rho_{\mathcal{B}}+\rho_{a}}\ ,$ (7)
where $M_{\rm Pl}\equiv(8\pi G)^{-1/2}\simeq 2.4\times 10^{18}\,{\rm GeV}$ is the reduced Planck mass.
@@ III.2 Axion temperature
Here, $\zeta(x)$ is the Riemann function for the argument $x$, and the normalization factor ensures that we recover the temperature of a Bose-Einstein distribution when the interactions are strong enough to ensure thermalization.
The definition in Eq. (8) is not the only one that reproduces the temperature of a Bose-Einstein. An analogous definition with the first moment of the distribution instead of the second would recover the equilibrium result upon choosing the normalization appropriately. However, for the most general distribution, the two results would differ. The scenario analyzed here, when equilibrium is not achieved, is analogous to dark matter freeze-in for which explicit distributions were derived in Ref. [49]. It was also shown that the temperatures defined via the first and second moments are quite similar. The only observable effect is $\Delta N_{\rm eff}$, the temperature is not observable no matter what definition we choose, we define a temperature is to have an estimate of the width of the distribution when we are far from equilibrium. In what follows, we will adopt the definition in Eq. (8) and show how the axion temperature evolves with time.
@@ IV Leptophilic Axions
If one chooses this clever set of variables then it is possible to investigate compactly in a single plot all deviations from thermal shapes for different choices of the coupling strengths. This is what the left panels of Fig. 2 illustrate. Axions produced from scatterings have a momentum distribution very close to a thermal shape identified by the black dotted line. This is true for large couplings where thermalization is achieved and for small couplings where axions do not get even close to equilibrium. The fact that the asymptotic distribution is always rather thermal up to an overall constant (the factor $R_{a}(A_{F})$ in the normalization) makes sense given the thermal origin. Crucially, the resulting temperatures are different as shown by the right panels of Fig. 2. Larger values of $f_{a}$ lead to smaller temperatures, and the net effect is a reduced contribution to $\Delta N_{\rm eff}$ that we evaluate in Sec. VI.
@@ Acknowledgements.
The authors thank Miguel Escudero, Rotem Ovadia and Luca Vecchi for useful discussions. We also thank the journal referee for their constructive feedback on the first version of this manuscript. This work is supported in part by the Italian MUR Departments of Excellence grant 2023-2027 “Quantum Frontiers”. F.D. is supported by Istituto Nazionale di Fisica Nucleare (INFN) through the Theoretical Astroparticle Physics (TAsP) project. F.D. acknowledges support from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 860881-HIDDeN. The work of A.L. is supported by an ERC STG grant (“Light-Dark”, grant No. 101040019). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement No. 101040019). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union. The European Union cannot be held responsible for them. A.L. is also in part supported by the US-Israeli BSF grant 2016153. A.L. is grateful to the CERN theory group for the support while this work was carried out.
@@ Appendix B Tracking the axion number density
$n_{a}(t)\equiv g_{a}\int\frac{d^{3}k}{(2\pi)^{3}}{\mathcal{F}}_{a}(k,t)\ .$ (20)